OpenStudy (anonymous):

–4n + 7 + 2n = 1 A) 1 B) 3 C) –3 D) 4

hartnn (hartnn):

combine like terms what is -4n+2n = ?

hartnn (hartnn):

?

OpenStudy (anonymous):

-2n @hartnn

OpenStudy (anonymous):

–4n + 7 + 2n = 1 Combine like terms. - 4n + 2n -2n + 7 = 1 Subtract 7 from both sides. - 7 -7 --------------- -2n = -6 Divide both sides by -2 /-2 /-2 --------------- n = 3

OpenStudy (anonymous):

oh that was easier than I thought or made it i just wondered about two variables